# MalPurifier: Enhancing Android Malware Detection with Adversarial Purification against Evasion Attacks

### v1

@@ 5       Flip feature values via $x_{i}[m]=1-x_{i}[m]$;
4.4 Accurate Sample Recovery
In this section, we introduce the technical details of the proposed purification mechanism and explain how to effectively recover perturbed adversarial malware samples while not affecting the detection of clean data through this mechanism. Our purification model is built on a denoising autoencoder, which has excellent capability to remove noise and improve the quality of the data for downstream tasks.
As illustrated in Fig. 2, the input of the purification model $\psi_{\vartheta}$ is the perturbed sample generated from diversified adversarial perturbation in Algorithm 1 (see Step 3), the modified benign samples created from protective noise injection in Algorithm 2 (see Step 4), as well as their original forms. During the training process of the purification model, these data are passed through and the network parameters $\vartheta$ can be optimized to minimize the customized loss that is described below.
Reconstruction Loss. A commonly used loss function for DAEs is the MSE loss, which measures the average squared difference between two groups of data. To help train a more robust purification model, we aim to minimize the discrepancy of the feature representation between the reconstructed output and the target data (i.e., reconstruction loss) as follows.
@@ 4.4 Accurate Sample Recovery
Prediction Loss. Given that the ultimate objective of adversarial purification is to enhance the classification performance of the subsequent malware detector, we propose using prediction loss to further optimize the performance of the purification model as follows.
@@
where $\Delta$ is formally defined in Eq. (9), and the distance measure used to compute $\Delta$ is MSE as well. In this way, the purification model will be strengthened in which the purified data will be closer to the original version in the internal representation of the prediction model (i.e., the malware detector), leading to more accurate prediction results.
Therefore, the overall loss function for $\psi_{\vartheta}$ is the combination of reconstruction loss and prediction loss as follows.
@@
where $\alpha$ and $\beta$ are the weights of two loss terms, and we have $\alpha+\beta=1$. The parameters $\vartheta$ will be optimized by minimizing the customized loss of the purification model.
@@ 5 Experiments and Evaluation
  • RQ2: Robustness against obfuscation attacks. How is the robustness of MalPurifier against obfuscation attacks?
  • RQ3: Robustness against oblivious attacks. How robust is MalPurifier against oblivious attacks where the attacker is unaware of the additional defense mechanism (e.g., the adversarial purifier $g$)?
  • RQ4: Robustness against adaptive attacks. How robust is MalPurifier against adaptive attacks in which the adversary has full knowledge of all defense mechanisms?
  • RQ5: Enhancement brought to other detectors. Can the purification model act as a plug to enhance other detectors against evasion attacks?
@@
  • ${}^{+}$ [42]. It enhances the robustness of the detector by another detector trained with an additional outlier class for detecting adversarial examples.
@@
  • FD-VAE [23]. It improves the DNN model by introducing an additional Variational AutoEncoder (VAE) for Feature Disentangle (FD) in different classes and combining their detection outcomes to make the final decision (FD-VAE).
@@
Metrics. The effectiveness of defenses is assessed using five standard metrics: False Positive Rate (FPR), False-Negative Rate (FNR), Accuracy (Acc), balanced Accuracy (bAcc), and F1 score. The inclusion of balanced Accuracy and F1 score accounts for the imbalanced nature of the dataset. In addition, we include training time to evaluate the overhead of these methods.
@@ 5.1 RQ1: Effectiveness and Cost without Attacks
[…]
@@
Results. Table I exhibits the effectiveness and overhead of all detectors when there is no attack. We observe that ${}^{+}$ achieves the highest detection accuracy (98.72% on Drebin and 99.23% on Androzoo) and F1 score (93.58% on Drebin and 99.56% on Androzoo), which are a little higher than those of the basic DNN model. The reason may be that adversarial training introduces extra adversarial examples that help ${}^{+}$ identify more malicious samples, resulting in a lower FNR and higher F1 score.
[…]
@@
Figure 3: The accuracy of different detectors against obfuscation attacks on Drebin and Androzoo datasets. The color gradient ranging from light to dark represents the increasing accuracy from low to high, with the effectiveness of MalPurifier against each attack annotated in the square.
5.2 RQ2: Robustness against Obfuscation Attacks
Experimental Setup. We measure the accuracy of all aforementioned methods under obfuscation attacks via the Drebin and Androzoo datasets. Since these attacks belong to black-box attacks and can be conducted in the problem space, therefore, we leverage obfuscation technology to modify malicious APKs and extract features from the modified versions. Specifically, we utilize an obfuscator called AVPASS [33], to wage 8 obfuscation-based attacks to perturb malware examples on the test set.
@@ 5.2 RQ2: Robustness against Obfuscation Attacks
Results. Fig. 3 illustrates the accuracy of the detectors on Drebin (left panel) and Androzoo (right panel) datasets under 8 obfuscation attacks. We make the first observation that DNN can not defeat all these attacks (accuracy $\leqslant$ 0.344% on Drebin and $\leqslant$ 5.871% on Androzoo), demonstrating that such attacks can hide malicious features to evade detection. Nevertheless, obfuscation attackers produce adversarial examples in a black-box manner, they cannot effectively evade these detectors except for the DNN model.
[…]
@@
Answer to RQ2: MalPurifier outperforms the other methods against all obfuscation attacks on the Drebin dataset (accuracy of 100%), and achieves accuracy $\geqslant$ 95% against 7 obfuscation attacks on the Androzoo dataset.
5.3 RQ3: Robustness against Oblivious Attacks
@@
Figure 4: The accuracy of different detectors against gradient-based oblivious attacks on Drebin (top panel) and Androzoo (bottom panel) datasets, along with the iteration ranging from 0 to 100.
[…]
@@
We separately wage 12 oblivious attacks on test malware examples. For BCA [36], Grosse [37], and PGD-$\ell_{1}$ [35] attacks, we perturb one feature per time with a maximum iteration of 100. For rFGSM [19] and $\ell_{\infty}$ [35], we iterate these attack algorithms with the iterations 100 and step size 0.02. The PGD-$\ell_{2}$ [35] attack is set with 100 iterations and step size 0.5. Moreover, we perform Salt & Pepper attack [13] by increasing the noise intensity of 0.001 each time until misclassification and repeating this process 10 times, and the Pointwise [38] attack utilizes Salt & Pepper as the initial attack and minimizes the needed perturbations.
@@
Additionally, we combine PGD-$\ell_{1}$, PGD-$\ell_{2}$, and $\ell_{\infty}$ attacks to perform MaxMA [20] attack and run it 5 times with the random starting point for the iMaxMA [20] attack. We also iterate the SMA [20] attack 100 times with the step size 0.5 for PGD-$\ell_{2}$ and 0.02 for $\ell_{\infty}$, and the AutoAttack [39] consists of APGD-CE and FAB with $\ell_{2}$ norm in our experiments.
@@
Results. Fig. 4 depicts the accuracy curves of these methods on Drebin (top panel) and Androzoo (bottom panel) datasets under 6 gradient-based oblivious attacks with the iteration from 0 to 100. We first observe an important observation that none of these attacks can evade MalPurifier (accuracy $\leqslant$ 90.91% on Drebin and $\leqslant$ 99.41% on Androzoo), demonstrating the high robustness of the proposed approach.
@@
We further observe that there is a decreasing trend on the curves of ${}^{+}$, KDE, and PAD-SMA against BCA, Grosse, and PGD-$\ell_{1}$ attacks until 20 iterations. This is because such attacks in an oblivious manner will stop manipulations when the adversarial example can evade the malware detector. Moreover, there exists a dip in some accuracy curves of ${}^{+}$, KDE, and FD-VAE against rFGSM and PGD-$\ell_{2}$ attacks with the iteration ranging from 0 to 100. This is because such attacks create small perturbations that can successfully escape on a specific iteration, and as the number of iterations increases, they will be recognized due to the larger perturbations.
@@
TABLE II: Accuracy of different defenses against oblivious attacks in which attackers only know the original malware detector.
@@
Answer to RQ3: MalPurifier is significantly more robust than ${}^{+}$, KDE, FD-VAE, and PAD-SMA. It achieves an accuracy of $\geqslant$ 90.91% and $\geqslant$ 99.24% on both datasets against oblivious attacks, wherein all of these adversarial examples are previously unseen by the model.
TABLE III: Accuracy of different defenses under adaptive attacks where adversaries know all defensive mechanisms (if applicable).
@@
5.4 RQ4: Robustness against Adaptive Attacks
[…]
@@ 5.4 RQ4: Robustness against Adaptive Attacks
Results. Table III reports the experimental results against adaptive attacks. First, we can observe that DNN is very vulnerable to all attacks, with 0% accuracy against 11 attacks on Drebin and 8% accuracy against 8 attacks on Androzoo. However, the Salt &\&& Pepper and Pointwise attacks achieve the lowest attack effectiveness in evading DNN on Androzoo, because both of them modify malware examples without using the internal information of target detectors.
[…]
@@
Third, although ${}^{+}$, KDE, and FD-VAE incorporate another adversary detector, they only show limited effectiveness under a few attacks, and suffer from unseen attacks such as PGD-$\ell_{1}$, MaxMA, iMaxMA, Orth PGD-$\ell_{1}$, Orth MaxMA, and Orth iMaxMA (with an accuracy $\leqslant$ 13.08% on both datasets). Additionally, PAD-SMA not only hardens the DNN with adversarial training, but also combines it with an ICNN-based adversary detector, significantly improving its robustness against different attacks. Especially, PAD-SMA achieves the highest accuracy of 98.34% against the PGD-$\ell_{2}$ attack on Androzoo. However, PAD-SMA is still sensitive to the Grosse, PGD-$\ell_{1}$, MaxMA, and iMaxMA attacks (with an accuracy $\leqslant$ 78.76% on Drebin and $\leqslant$ 56.30% on Androzoo). Considering its high time overhead reported in Table I, it cannot be called a perfect solution.
@@
In summary, MalPurifier significantly outperforms other defenses, achieving the highest accuracy against all 17 adaptive attacks on the Drebin dataset and 15 attacks on the Androzoo dataset. This indicates that the purification model can accurately recover the original forms of adversarial examples even if the adversary knows its existence.
@@
Answer to RQ4: MalPurifier outperforms other defenses in the condition that adversaries know all defensive mechanisms, and significantly hardens the malware detector against a wide range of adaptive attacks (with an accuracy $\geqslant$ 90.91% on Drebin and $\geqslant$ 97.44% on Androzoo).
5.5 RQ5: Enhancement Brought to Other Detectors

### v2

@@ 5       Flip feature values via $x_{i}[m]=1-x_{i}[m]$;
4.4 DAE-based Purification Model
Building on the mechanisms described in Sections 4.2 and 4.3, we construct a comprehensive training dataset by combining diversified adversarial malware samples, benign samples with protective noise, and their original clean counterparts. Therefore, in this section, we present the technical details of our DAE-based purification model and explain how it effectively restores perturbed malware samples while preserving the detection accuracy for clean data.
While various generative models can be used for purification, the choice of model is critical for handling the unique nature of malware features. We specifically adopt a DAE for the Android malware detection. This choice is motivated by three main reasons: (i) Compared to classical methods like Principal Component Analysis (PCA), the DAE is particularly effective at removing diverse and structured noise from discrete, high-dimensional feature spaces, which matches the binary and sparse nature of Android malware features. (ii) Unlike generative models such as Variational Autoencoder (VAE) and Generative Adversarial Network (GAN), the DAE does not require adversarial objectives or complex regularization, making it more scalable and robust in practice. (iii) Compared to Diffusion Model (DM), the DAE is easier to train and model-agnostic, allowing it to be seamlessly integrated as a plug-and-play purification module for various downstream detectors.
As illustrated in Fig. 2, the purification model $\psi_{\vartheta}$ is trained with several types of input: perturbed malware samples (generated via Algorithm 1), benign samples modified by protective noise (via Algorithm 2), as well as their original, unperturbed forms. During training, these data are passed through the network, and the parameters $\vartheta$ are optimized to minimize a customized loss function as described below.
Reconstruction Loss. A standard loss function for DAE is the MSE loss, which measures the average squared difference between the reconstructed output and the target data. To train a robust purification model, we minimize the discrepancy between the reconstructed output and the original data (i.e., reconstruction loss) as follows.
@@ 4.4 DAE-based Purification Model
Prediction Loss. Since the ultimate goal of adversarial purification is to improve the classification performance of the downstream malware detector, we introduce a prediction loss to further optimize the purification model as follows.
@@
where $\Delta$ is formally defined in Eq. (9), and $d(\cdot)$ is again the MSE. This loss encourages the purified data to be close to the original data in the internal feature space of the malware detector, leading to more accurate predictions.
The overall loss function for training $\psi_{\vartheta}$ is a weighted combination of the reconstruction loss and the prediction loss as follows.
@@
where $\alpha$ and $\beta$ are the weights of two loss terms, and we have $\alpha+\beta=1$. The parameters $\vartheta$ are optimized by minimizing this combined loss during training.
@@ 5 Experiments and Evaluation
  • RQ2: Robustness against black-box attacks. How is the robustness of MalPurifier against black-box attacks?
  • RQ3: Robustness against grey-box attacks. How robust is MalPurifier against grey-box attacks where the attacker is unaware of the additional defense mechanism (e.g., the adversarial purifier $g$)?
  • RQ4: Robustness against white-box attacks. How robust is MalPurifier against white-box attacks in which the adversary has full knowledge of all defense mechanisms?
  • RQ5: Advantage and transferability of the purifier. Does the DAE model in MalPurifier outperform alternative purification techniques, and can it be flexibly transferred to enhance other types of detectors against evasion attacks?
  • RQ6: Generalizability against advanced attacks. How does MalPurifier perform against advanced attacks that target the structural or semantic properties of Android applications?
@@
  • DNN+ [42]. It enhances the robustness of the detector by another detector trained with an additional outlier class for detecting adversarial examples.
@@
  • FD-VAE [22]. It improves the DNN model by introducing an additional VAE for Feature Disentangle (FD) in different classes and combining their detection outcomes to make the final decision (FD-VAE).
@@
Metrics. The effectiveness of defenses is assessed using five standard metrics as follows. False Positive Rate (FPR) denotes the proportion of benign samples incorrectly classified as malicious and False Negative Rate (FNR) represents the proportion of malicious samples incorrectly classified as benign. Accuracy (Acc) is the percentage of the test examples that are correctly while balanced Accuracy (bAcc) can be defined as the average accuracy obtained on either class. F1 score denotes a harmonic mean of precision and recall that combines the performance of a classifier in terms of both false positives and false negatives. In addition, we include training time to evaluate the overhead of these methods.
@@ 5.1 RQ1: Effectiveness and Cost without Attacks
In detail, DNN+ [42] leverages another detector hardened by adversarial training with the MaxMA attack against the DNN model to identify adversarial examples, and KDE [43] relies on the close distance between activations to reject large manipulations without retraining. FD-VAE [22] incorporates a VAE-based indicator to classify clean data and adversarial examples, in which both the encoder and decoder consist of two layers (each layer having 600 neurons) with the Softplus activation. Moreover, AT-rFGSMk [35] uses the $\ell_{\infty}$ attack, which has 50 iterations with step size 0.02, and AT-Adam [44] exploits the Adam optimizer with iterations 50, step size 0.02, and random starting point. PAD-SMA [9] uses three attacks, including PGD-$\ell_{1}$ attack iterates 50 times, PGD-$\ell_{2}$ attack iterates 50 times with step size 0.5, and $\ell_{\infty}$ iterates 50 times with step size 0.02.
@@
Results. Table I exhibits the effectiveness and overhead of all detectors when there is no attack. We observe that DNN+ achieves the highest detection accuracy (98.72% on Drebin and 99.23% on Androzoo) and F1 score (93.58% on Drebin and 99.56% on Androzoo), which are a little higher than those of the basic DNN model. The reason may be that adversarial training introduces extra adversarial examples that help DNN+ identify more malicious samples, resulting in a lower FNR and higher F1 score.
An interesting observation is that KDE takes the same time as DNN whereas the other methods have higher overhead in the training phase. The underlying reason may be that the KDE method builds another KDE-based detector without retraining, while the others train a separative model to detect adversarial examples upon the DNN model. We further observe that MalPurifier’s FNR decreases but FPR increases, leading to decreased accuracy (1.72% on Drebin and 0.66% on Androzoo) on clean data. which is similar to that of adversarial training methods (e.g., AT-rFGSMk, AT-Adam, and PAD-SMA). This is because our focus primarily lies on ”purifying” adversarial examples into original forms that can be detected by the following DNN-based detector, although we do process benign samples as well.
@@
Figure 3: The accuracy of different detectors against black-box attacks on Drebin and Androzoo datasets. The color gradient ranging from light to dark represents the increasing accuracy from low to high, with the effectiveness of MalPurifier against each attack annotated in the square.
5.2 RQ2: Robustness against Black-Box Attacks
Experimental Setup. After establishing MalPurifier’s baseline performance on clean data, we now investigate its effectiveness against black-box attacks. In detail, we measure the accuracy of all aforementioned methods under obfuscation attacks. These attacks utilize obfuscation technology to modify and conceal malicious functionality without the knowledge of the target classifier. Specifically, we utilize an obfuscator called AVPASS [32], to wage 8 kinds of attacks to perturb malware examples and extract features from the modified versions on the test set.
@@ 5.2 RQ2: Robustness against Black-Box Attacks
Results. Fig. 3 illustrates the accuracy of the detectors on Drebin (left panel) and Androzoo (right panel) datasets under 8 obfuscation-based black-box attacks. We make the first observation that DNN can not defeat all these attacks (accuracy $\leqslant$ 0.344% on Drebin and $\leqslant$ 5.871% on Androzoo), demonstrating that such attacks can hide malicious features to evade detection. Nevertheless, attackers produce adversarial examples in a black-box manner, they cannot effectively evade these detectors except for the DNN model.
We further observe significant differences between the robustness of some detectors against these attacks on the two datasets. For example, DNN+ shows poor performance on Drebin whereas achieves high robustness on Androzoo. This can be attributed to the fact that the sample structures used in the two datasets are significantly different (Drebin collected in 2013 whereas Androzoo collected in 2021) and the data imbalance may lead to this phenomenon as well. Another interesting observation is that some adversarial training methods (e.g., AT-rFGSMk, AT-Adam) show high robustness on black-box attacks, which may be attributed to the similarity between adversarial examples generated by PGD attacks and obfuscation technology.
@@
Answer to RQ2: MalPurifier outperforms the other methods against all black-box attacks on the Drebin dataset (accuracy of 100%), and achieves accuracy $\geqslant$ 95% against 7 black-box attacks on the Androzoo dataset.
5.3 RQ3: Robustness against Grey-Box Attacks
@@
Figure 4: The accuracy of different detectors against gradient-based grey-box attacks on Drebin (top panel) and Androzoo (bottom panel) datasets, along with the iteration ranging from 0 to 100.
Experimental Setup. Having demonstrated MalPurifier’s resilience against black-box attacks, we now evaluate its performance against more challenging grey-box attacks. In these attacks, the adversaries are aware of the baseline DNN classifier but remain oblivious to other defensive mechanisms. Since DNN, AT-rFGSMk, and AT-Adam lack additional detectors or indicators, we solely consider the robustness of DNN+, KDE, FD-VAE, PAD-SMA, and MalPurifier against 12 grey-box attacks.
@@
First, we wage 6 gradient-based grey-box attacks in an oblivious manner on test malware examples. For BCA [35], Grosse [36], and PGD-$\ell_{1}$ [34] attacks, we perturb one feature per time with a maximum 100 iterations. For rFGSM [18] and $\ell_{\infty}$ [34], we iterate these attack algorithms with 100 iterations and a step size of 0.02. The PGD-$\ell_{2}$ [34] attack is set with 100 iterations and a step size of 0.5.
We also incorporate 2 gradient-free attacks in the grey-box scenario. We conduct Salt & Pepper attack [12] by increasing the noise intensity of 0.001 each time until misclassification and repeating this process 10 times. The Pointwise [37] attack utilizes Salt & Pepper as the initial attack and minimizes the needed perturbations.
@@
Furthermore, 4 ensemble-based grey-box attacks are included in this section. We combine PGD-$\ell_{1}$, PGD-$\ell_{2}$, and $\ell_{\infty}$ to perform MaxMA [19] attack and run it 5 times with the random starting point for the iMaxMA [19] attack. We iterate the SMA [19] attack 100 times with a step size of 0.5 for PGD-$\ell_{2}$ and 0.02 for $\ell_{\infty}$, and the AutoAttack [38] comprises APGD-CE and FAB with the $\ell_{2}$ norm.
@@
Results. Fig. 4 depicts the accuracy curves of these methods on Drebin (top panel) and Androzoo (bottom panel) datasets under 6 gradient-based grey-box attacks with the iteration from 0 to 100. We first observe an important observation that none of these attacks can evade MalPurifier (accuracy $\leqslant$ 90.91% on Drebin and $\leqslant$ 99.41% on Androzoo), demonstrating the high robustness of the proposed approach.
@@
We further observe that there is a decreasing trend on the curves of DNN+, KDE, and PAD-SMA against BCA, Grosse, and PGD-$\ell_{1}$ attacks until 20 iterations. This is because such attacks in an oblivious manner will stop manipulations when the adversarial example can evade the malware detector. Moreover, there exists a dip in some accuracy curves of DNN+, KDE, and FD-VAE against rFGSM and PGD-$\ell_{2}$ attacks with the iteration ranging from 0 to 100. This is because such attacks create small perturbations that can successfully escape on a specific iteration, and as the number of iterations increases, they will be recognized due to the larger perturbations.
@@
TABLE II: Accuracy of different defenses against gradient-free and ensemble-based grey-box attacks on Drebin and Androzoo datasets.
@@
Answer to RQ3: MalPurifier is significantly more robust than DNN+, KDE, FD-VAE, and PAD-SMA. It achieves an accuracy of $\geqslant$ 90.91% and $\geqslant$ 99.24% on both datasets against grey-box attacks, wherein all of these adversarial examples are previously unseen by the model.
TABLE III: Accuracy of different defenses under white-box attacks where adversaries know all defensive mechanisms (if applicable).
@@
5.4 RQ4: Robustness against White-Box Attacks
Experimental Setup. To further evaluate MalPurifier under worst-case conditions, we now progress to white-box attacks, where attackers have complete knowledge of both the detector $f$ and the adversarial indicator or purifier $g$.
First, we adapt the 12 grey-box attacks to white-box attacks by solving the problem in Eq. 7. Since DNN, AT-rFGSMk, and AT-Adam do not have any adversarial indicator, the grey-box attacks aimed at these defenses can trivially fulfill the adaptive requirement of white-box attacks.
Furthermore, we improve the other 5 white-box attacks by producing perturbations into two components (e.g., detector $f$ and purifier $g$) in an ”orthogonal” (dubbed Orth) manner [45] to prevent perturbation waste, including Orth PGD-$\ell_{1}$, PGD-$\ell_{2}$, $\ell_{\infty}$, MaxMA, and iMaxMA. For similar reasons, these attacks are not applicable to DNN, AT-rFGSMk, and AT-Adam approaches. Note that we utilize the same hyper-parameters as those in Section 5.3, except for PGD-$\ell_{1}$ with 500 iterations, PGD-$\ell_{2}$ with 200 iterations and step size 0.05, and $\ell_{\infty}$ with 500 iterations and step size 0.002 in all PGD-based attacks.
@@ 5.4 RQ4: Robustness against White-Box Attacks
Results. Table III reports the experimental results against white-box attacks. First, we can observe that DNN is very vulnerable to all attacks, with 0% accuracy against 11 attacks on Drebin and 8% accuracy against 8 attacks on Androzoo. However, the Salt &\&& Pepper and Pointwise attacks achieve the lowest attack effectiveness in evading DNN on Androzoo, because both of them modify malware examples without using the internal information of target detectors.
Second, adversarial training methods (e.g., AT-rFGSMk, and AT-Adam) can harden the robustness of DNN to some extent. For example, AT-rFGSMk can mitigate the Salt &\&& Pepper, Pointwise, and AutoAttack attacks, and AT-Adam is also effective in defeating rFGSM, PGD-$\ell_{2}$ and $\ell_{\infty}$ attacks. Nevertheless, they are still sensitive to BCA, Grosse, PGD-$\ell_{1}$, MaxMA, iMaxMA, and StepwiseMA attacks (with an accuracy $\leqslant$ 45.83% on both datasets) that are unseen previously. These results indicate the limitations of adversarial training methods in terms of generalization.
@@
Third, although DNN+, KDE, and FD-VAE incorporate another adversary detector, they only show limited effectiveness under a few attacks, and suffer from unseen attacks such as PGD-$\ell_{1}$, MaxMA, iMaxMA, Orth PGD-$\ell_{1}$, Orth MaxMA, and Orth iMaxMA (with an accuracy $\leqslant$ 13.08% on both datasets). Additionally, PAD-SMA not only hardens the DNN with adversarial training, but also combines it with an ICNN-based adversary detector, significantly improving its robustness against different attacks. Especially, PAD-SMA achieves the highest accuracy of 98.34% against the PGD-$\ell_{2}$ attack on Androzoo. However, PAD-SMA is still sensitive to the Grosse, PGD-$\ell_{1}$, MaxMA, and iMaxMA attacks (with an accuracy $\leqslant$ 78.76% on Drebin and $\leqslant$ 56.30% on Androzoo). Considering its high time overhead reported in Table I, it cannot be called a perfect solution.
@@
In summary, MalPurifier significantly outperforms other defenses, achieving the highest accuracy against all 17 white-box attacks on the Drebin dataset and 15 attacks on the Androzoo dataset. This indicates that the purification model can accurately recover the original forms of adversarial examples even if the adversary knows its existence.
@@
Answer to RQ4: MalPurifier outperforms other defenses in the condition that adversaries know all defensive mechanisms, and significantly hardens the malware detector against a wide range of white-box attacks (with an accuracy $\geqslant$ 90.91% on Drebin and $\geqslant$ 97.44% on Androzoo).
5.5 RQ5: Advantage and Transferability of the Purifier
Experimental Setup. While previous experiments have demonstrated MalPurifier’s effectiveness against increasingly sophisticated attacks, we now investigate two critical aspects: (i) whether our choice of DAE as the purification model is optimal compared to alternative techniques, and (ii) whether MalPurifier can be flexibly integrated with different detection architectures.
First, to justify the effectiveness of the DAE-based purification model, we compare it with alternative purification techniques, including VAE, PCA, GAN, and DM. The DAE model is trained with the same hyper-parameters as that in Section 5.1. VAE shares the same architecture of the encoder and decoder as DAE, but with the Softplus activation, and it introduces a Kullback–Leibler (KL) weight of 1.0 and regularization coefficient of 0.001. For PCA, the number of principal components is selected to retain more than 95% of the variance, and the projection matrix is derived from the training data. Both generator and discriminator of the GAN model use a hidden size of 600, and the latent dimension is set to 256 with a gradient penalty of 10. DM uses a sinusoidal time embedding with a dimension of 256, and a two-layer Multilayer Perceptron (MLP) with a hidden size of 600 to predict noise residuals at arbitrary steps.
For comprehensive comparison among different purification methods, we select several test scenarios including a clean setting and five white-box attacks: rFGSM, $\ell_{\infty}$, Salt & Pepper, MaxMA, and StepwiseMA. Attack configurations follow the setup described in Section 5.4, and we use a DNN-based model as the classifier with the same hyper-parameters as that in Section 5.1.
Second, to further evaluate the flexibility and transferability of the purification model, we also package the DAE model as a plug to protect other detectors, such as Support Vector Machine (SVM), Fully Convolutional Network (FCN), Long Short Term Memory (LSTM), and Recurrent Neural Network (RNN). In detail, We use an SVM model with the Sigmoid activation and an FCN model with 3 fully-connected hidden layers (each having 512, 256, and 128 neurons) with the ReLU activation. The LSTM model has a hidden layer with 200 neurons, a sequence length of 1, and uses the Sigmoid function as the activation. Furthermore, we build the RNN model with 3 hidden layers (each having 200 neurons) and the Sigmoid activation. All these models are tuned by the Adam optimizer with 100 epochs, batch size 128, and learning rate 0.001.
Note that the DAE-based purification model works as a plug-and-play preprocessing method, in which we do not retrain the model but directly apply it from the aforementioned experiments. In addition, we wage 6 attacks on test malware examples in a white-box manner, including BCA, PGD-$\ell_{1}$, PGD-$\ell_{2}$, $\ell_{\infty}$, Pointwise, and StepwiseMA, with the same hyper-parameters in Section 5.4.
Figure 5: The accuracy of different purification algorithms when equipped with the DNN-based classifier in the absence and presence of evasion attacks on Drebin and Androzoo datasets.